Solve each equation

(a) p^2-8p=0

(b) (x-3)(x+5)=-7 (c) (3x+2)(x-3)=7x-1

Answer by stanbon(75887)

Solve each equation
(a) p^2-8p=0
p(p-8) = 0
p = 0 or p = 8
----------------------
(b) (x-3)(x+5)=-7
x^2+2x-15 = -7
x^2+2x-8 = 0
(x+4)(x-2) = 0
x = -4 or x = 2
----------------------
(c) (3x+2)(x-3)=7x-1
3x^2-7x-6 = 7x-1
3x^2-14x-5 = 0
---
3x^2-15x+x-5 = 0
3x(x-5)+(x-5) = 0
(x-5)(3x+1) = 0
x = 5 or x = -1/3
=====================
Cheers,
Stan H.
===========